# Mapping Trustworthiness in Large Language Models: A Bibliometric Analysis Bridging Theory to Practice

### v1

@@
Abstract.
The rapid proliferation of Large Language Models (LLMs) has raised pressing concerns regarding their trustworthiness, spanning issues of reliability, transparency, fairness, and ethical alignment. Despite the increasing adoption of LLMs across various domains, there remains a lack of consensus on how to operationalize trustworthiness in practice. This study bridges the gap between theoretical discussions and implementation by conducting a bibliometric mapping analysis of 2,006 publications from 2019 to 2025. Through co-authorship networks, keyword co-occurrence analysis, and thematic evolution tracking, we identify key research trends, influential authors, and prevailing definitions of LLM trustworthiness. Additionally, a systematic review of 68 core papers is conducted to examine conceptualizations of trust and their practical implications. Our findings reveal that trustworthiness in LLMs is often framed through existing organizational trust frameworks, emphasizing dimensions such as ability, benevolence, and integrity. However, a significant gap exists in translating these principles into concrete development strategies. To address this, we propose a structured mapping of 20 trust-enhancing techniques across the LLM lifecycle, including retrieval-augmented generation (RAG), explainability techniques, and post-training audits. By synthesizing bibliometric insights with practical strategies, this study contributes towards fostering more transparent, accountable, and ethically aligned LLMs, ensuring their responsible deployment in real-world applications.
@@ Abstract.
copyright: acmlicensedjournalyear: 1014doi: XXXXXXX.XXXXXXXconference: The Xth International Conference on XXXX; XX–$X$ June, 2025; XXXX, XXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: General and reference Surveys and overviewsccs: Software and its engineering Software development techniquesccs: Social and professional topics Computing / technology policy
1. Introduction
@@
Artificial Intelligence (AI) has become a cornerstone of technological progress, yet remains at the centre of ethical debate. More than 200 guidelines and principles have been proposed to guide its development and use, reflecting the growing recognition of its societal impact (Hagendorff, 2020; Corrêa et al., 2023). These AI ethical guidelines, which have been proposed by public organisations, society, academia and private companies, do not provide objective guidance to developers, but rather abstract principles that are not practical for developers to operationalise, e.g, transparency, privacy, justice & fairness (Vakkuri et al., 2021; Floridi and Cowls, 2022; de Cerqueira et al., 2022).
@@
As an evolution of AI system that employs advanced deep learning techniques, Large Language Models (LLMs) are trained on vast datasets and capable of generating novel text, images, audio, and video, based on the data it was trained on (Sun et al., 2024). Their rapid adoption has sparked both enthusiasm and concern, as their potential for negative outcomes – such as spreading misinformation, perpetuating biases, or hallucination – has become evident (Liu et al., 2023; Wang et al., 2023). Despite efforts like the Paris AI Action Summit (Élysée, 2025), which seeks to promote responsible AI development and international cooperation, significant challenges remain. For instance, Google’s recent decision to remove its AI principles prohibiting the use of AI in autonomous weapons and surveillance has raised questions about the commitment of major tech companies to ethical AI (Hooker and Vallance, 2025). This shift underscores the tension between ethical ideals and commercial or geopolitical interests, highlighting the need for robust, enforceable legislations.
@@
The swift popularization of LLMs have encouraged researchers in understanding the specific ethical dimensions of this new generative AI (genAI), introducing them mainly as trustworthiness taxonomies, i.e., a rebranded set of similar abstract AI ethical principles (Sun et al., 2024; Liu et al., 2023). Trustworthiness in LLMs can be characterized in terms of reliability, safety, fairness, resistance to misuse, explainability & reasoning, social norm and robustness (Liu et al., 2023). Thus, the focus of AI ethics discussions has shifted to the trustworthiness of LLM systems. The current debate on LLM trustworthiness largely focus on providing definitions, benchmarks and evaluation methods (Sun et al., 2024; Liu et al., 2023; Wang et al., 2023).
@@
Analogously with AI ethics, the proliferation of trustworthiness taxonomies for LLMs has not led to a clear consensus on definitions or a practical framework to guide the development of these systems (Smith et al., 2025). Despite the growing body of literature on LLM trustworthiness, there is still no unified set of practical guidelines to help developers mitigate risks throughout the development and deployment process. This gap between theory and practice not only hinders the creation of more trustworthy LLM systems but also masks deeper structural problems that have persisted since the early days of AI ethics.
This study aims to bridge the gap between theory and practice by investigating the means of operationalising trustworthiness in LLMs. We conduct a bibliometric mapping study, with 2,006 studies from 2019 to 2025, to identify the most influential authors, research trends and approaches and systematically analyse their similarities and differences. Additionally, by manually reading 68 key documents, we also conduct an analysis of the different definitions of trustworthiness proposed in the literature, providing insights into how different authors conceptualise trust in LLMs. Furthermore, we extend our analysis by identifying practical strategies that developers can use to enhance the trustworthiness of LLMs throughout the development lifecycle – from pre-training to deployment and real-world use. By bringing these perspectives together, we aim to provide both a theoretical and practical basis for promoting more ethical and trustworthy LLMs.
@@
  • Bibliometric Mapping Analysis: Using Bibliometrix, an R-tool, we conduct a bibliometric mapping study to identify the most influential authors, definitions, and approaches in the field. This analysis reveals trends, patterns, and areas of divergence in the literature on LLM trustworthiness.
  • Systematic Analysis of Trustworthiness Definitions: We compile and categorize various definitions of trust and trustworthiness in LLMs proposed by different authors. By analyzing these definitions, we identify commonalities, differences, and conceptual overlaps with related terms such as Responsible AI and AI Ethics.
  • LLM Trustworthiness Enhancing Strategies: We propose a mapping of practical techniques, technologies and processess that developers can use to enhance LLM trustworthiness at various stages of the development lifecycle, from pre-training to post-training, deployment and end-use. This helps in bridging the gap between abstract taxonomies and actionable steps.
  • Actionable Recommendations: Based on our findings, we offer actionable recommendations for developers, policymakers, and researchers to operationalize trustworthiness principles in practice. These recommendations aim to address the structural challenges that have hindered the development of trustworthy AI systems.
By integrating a systematic review of trustworthiness definitions with bibliometric insights and practical techniques, this work contributes to ongoing efforts to operationalize trustworthiness in LLMs and offers concrete steps toward the development of more trustworthy LLM systems.
This article is organized as follows: in the Section 2, we review similar works that conduct literature reviews on AI ethical guidelines and LLM trustworthiness taxonomies, highlighting their similarities and differences. In Section 3, we describe the Bibliometrix methodology adopted. In Section 4, we present the main findings, including systematic analyses of LLM trustworthiness definitions and trustworthiness-enhancing techniques. In Section LABEL:threats_to_validity, we discuss the study’s limitations and potential biases. Finally, in Section LABEL:final_remarks_section, we reflect on the practical and theoretical implications of the results, suggesting future directions for research and development.
2. Related Work
@@
In the context of this study, we examine previous literature reviews related to LLM and genAI trust/trustworthiness and ethics in the context of this study. Among the most relevant works is that of Liu et al. (Liu et al., 2023), which proposes a taxonomy for evaluating trustworthiness in LLM in seven major categeories: reliability, safety, fairness, resistance to misuse, explainability and reasoning, adherence to social norms, and robustness. Although the paper is not a literature review, it does describe each of the major categories with examples. However, the authors fail to provide a clear definition of what is trustworthy and do not provide practical guidance for developers on how to operationalise their taxonomy.
@@
Albahri et al. (Albahri et al., 2023) conducted a systematic science mapping analysis in the context of trustworthy and explainable AI (XAI) in healthcare systems, and found various XAI techniques for developers to operationalise. Unlike their approach, we aim to provide means to operationalise more of the trustworthiness categories.
It can be seen that many scholars have provided reviews of AI ethical principles, such as (Mittelstadt, 2019), (Jobin et al., 2019), (Corrêa et al., 2023), (Hagendorff, 2020), (Siqueira De Cerqueira et al., 2021), also providing an overview of publicly available tools that can help operationalise ethical principles in AI (Morley et al., 2019a). However, they are not as relevant to our review as they appeared before the LLM and genAI hype, and we want to explore the notion of trustworthiness in LLM and genAI and how it relates to ethics in AI, also providing practical guidance for developers to operationalise it.
3. Methodology
This study employs a bibliometric analysis to systematically investigate the landscape of research on trustworthiness in LLMs. Bibliometric analysis is a quantitative research method used to identify trends, influential works, and key research themes in a given domain (Donthu et al., 2021). To achieve this, we utilize Bibliometrix, an open-source R package specifically designed for bibliometric and scientometric studies (Aria and Cuccurullo, 2017).
3.1. Data Collection
@@
The dataset for this study was extracted from the Web of Science (WoS) Core Collection during 21st of February 2025, as this database contain high-quality, peer-reviewed publications in Artificial Intelligence, ethics, and Software Engineering, and provides the highest quality of publications and reliability in its indexing of high-ranking journals. (Caputo et al., 2021). The search query was formulated using a combination of relevant keywords to ensure a comprehensive dataset:
TS=((”Large Language Model*” OR LLM OR ”Generative AI” OR ”genAI”) AND (Trust* OR Ethic*))
@@
After filtering out duplicates, irrelevant papers, and LLM as authors abbreviation, a final dataset of 2,006 publications was obtained. This amount of studies discussing LLMs is impressive, since we had to reduce the year range of our search to only 6 years due to the fact that LLM is a new technology.
@@
Further manual reading: After exploring the broader set with Bibliometrix, we conducted a manual reading of a reduced dataset in order to gain insightful information and cross-validate our findings. This manual reading was undertaken to provide answers to research questions 2 and 3. This minimal set of studies was obtained by filtering within Bibliometrix. An objective way to define whether or not there was a focus on trustworthiness was to look at the discussion at the abstract level. Thus, by filtering papers that had ”trust*” in the abstract, a set of 68 papers was found for manual reading.
3.2. Data Processing and Preprocessing
The raw bibliographic data was exported in BibTeX format and processed using Bibliometrix in R. The following preprocessing steps were conducted:
  • Data Cleaning: Removal of missing entries, duplicate records, and irrelevant fields.
  • Normalization: Unification of author names and keyword variations to ensure consistency.
3.3. Bibliometric Analysis
The bibliometric analysis consisted of the following techniques:
  • Descriptive Analysis: Calculation of publication trends, citation statistics, and author productivity over time.
  • Co-Authorship Network Analysis: Mapping collaborations between authors and institutions to identify key contributors in LLM trustworthiness research.
  • Co-Citation and Bibliographic Coupling: Identifying clusters of highly interconnected research papers and their thematic evolution.
  • Keyword Co-Occurrence Analysis: Investigating the most frequently occurring terms to identify emerging topics and research gaps.
  • Thematic Mapping: Using Biblioshiny (a visualization tool within Bibliometrix) to classify research themes based on centrality and density.
3.4. Visualization and Interpretation
To facilitate a deeper understanding of the research trends, we generated various network visualizations, including:
  • Co-Authorship Networks (to highlight key collaborations).
  • Keyword Co-Occurrence Networks (to reveal thematic clusters).
  • Citation Networks (to track the evolution of influential studies).
The results of these analyses were synthesized to provide insights into the state-of-the-art, research gaps, and future directions in LLM trustworthiness research.
4. Results and Discussion
@@
This study examines the landscape of trustworthiness and ethics research in LLMs and generative AI from 2019 to 2025. A total of 2006 relevant studies, published across 1160 different venues over the past six years, were analyzed. These publications were authored by 7901 researchers, with an average citation count of 8.901 per document. Notably, collaborative research dominates the field, with 96.6% of the studies involving multiple authors (7,632 in total), while only 3.4% were produced by single authors (269 individuals).
4.1. What are the research trends, key authors, and thematic areas in the study of LLM trustworthiness?
4.2. Evolution of Publications from 2019-2025
@@
The study shows a sharp rise in publications on LLM trustworthiness and ethics from 2019 to early 2025, growing from 3 articles in 2019 to 1,434 in 2024, shown in Table 1 and in Figure 1. This growth was driven by advances like GPT-3/4, regulatory debates (e.g., EU AI Act), and societal concerns about AI ethics. Collaborative research dominated, with 96.6% of studies involving multiple authors. By February 2025, 213 articles were published, suggesting continued momentum ( 1,278 projected by year-end). This highlights the increase interest researchers have in trust/trustworthiness and ethical LLM and genAI research.
Table 1. Number of Articles Published Per Year
Year Articles
2019 3
2020 1
2021 4
2022 12
2023 339
2024 1434
2025 213
Figure 1. Evolution of Documents Per Year - 2019 - 2025.
4.3. Most Global Cited Documents
Figure 2 shows the top documents ranked by total number of citations, without any filtering. Articles with high global citations indicate significant impact and influence across disciplines (Donthu et al., 2021).
The document ”ChatGPT for good? On opportunities and challenges of large language models for education” by Kasneci et al. (Kasneci et al., 2023) has received 1,135 global citations. The paper explores the opportunities and challenges of using LLMs like ChatGPT in education, highlighting their potential to personalize learning, create educational content, and support both students and teachers. However, it also addresses critical risks such as bias, over-reliance, ethical concerns, and sustainability, emphasizing the need for responsible integration, teacher training, and ongoing research. The authors advocate for a balanced approach that leverages LLMs to enhance education while ensuring ethical use, transparency, and equitable access. Ongoing efforts to perceive the use of LLMs in education highlights how academia is concerned about this disruptive tool in educational settings.
The second most cited document is ”Opinion Paper: “So what if ChatGPT wrote it?” Multidisciplinary perspectives on opportunities, challenges and implications of generative conversational AI for research, practice and policy” by Dwivedi et al. (Dwivedi et al., 2023), with an equal amount of 1,135 global citations. The paper explores the multidisciplinary impact of ChatGPT, highlighting its potential to enhance productivity across industries while raising concerns about ethics, misinformation, and transparency. It discusses its role in research, business, and education, emphasizing both opportunities and challenges, including AI’s effect on academic integrity and hybrid work. The authors propose multiple future research directions, e.g., on AI governance, human-AI collaboration, and responsible implementation.
The third most cited document ”ChatGPT Utility in Healthcare Education, Research, and Practice: Systematic Review on the Promising Perspectives and Valid Concerns” by Sallam Malik (Sallam, 2023). This study systematically reviews the potential applications and limitations of ChatGPT in healthcare. It highlights the benefits of ChatGPT in scientific writing, research efficiency, medical education, and clinical workflows, while addressing concerns such as misinformation, bias, ethical and legal challenges, and safety risks. The study emphasises the need for responsible use of AI in healthcare and calls for regulatory guidelines and stakeholder collaboration to ensure ethical and effective implementation.
The three highly cited documents analysed present investigations of the specific adoption and impact of ChatGPT in different domains – education (Kasneci et al., 2023), research, business, and education (Dwivedi et al., 2023), and healthcare (Sallam, 2023). This shows how the research community has responded to the rapid emergence of ChatGPT in several domains. Furthermore, all three top documents have in common the need to ensure the ethical use, transparency and responsible implementation and use of AI.
Figure 2. Most Globally Cited Documents (2022-2024).
4.4. Most Local Cited Sources
When analyzing the most cited sources in our dataset in Table 2, we found an overwhelming 10,237 citations to arXiv preprint documents. This highlights a strong reliance on this preprint repository, suggesting that research in this area is rapidly evolving, with many studies referencing early-stage, non-peer-reviewed work. The prominence of arXiv, a widely used open-access platform for sharing preprints in fields like computer science, physics, and artificial intelligence, reflects the fast-paced nature of this research domain, where immediate access to emerging findings is valuable. However, while arXiV is interesting for the open-access, the documents inside are mostly pre-prints, that is, they mostly lack the scientific rigour that draws from the peer-review process.
In second place is ADV NEUR IN with 1,021 citations, which stands for Advances in Neural Information Processing Systems (NeurIPS). This is one of the most prestigious conferences in artificial intelligence and machine learning, attracting top researchers from around the world. It serves as a key platform for presenting groundbreaking advances in deep learning, reinforcement learning, and computational neuroscience that are shaping the future of AI research. The high citation rate of NeurIPS papers reflects its significant impact, as cutting-edge developments presented at the conference often drive innovation across academia and industry.
Nature comes in third with 980 citations. Nature is one of the world’s most prestigious and influential scientific journals, publishing ground-breaking research in a wide range of disciplines, including artificial intelligence, medicine and environmental science. Its rigorous peer-review process and high impact factor make it a trusted source for cutting-edge discoveries that shape scientific progress. The high citation rate of Nature articles reflects their far-reaching impact, often setting new research directions and informing policy and industry advances.
The presence of Journal of Medical Internet Research (JMIR) with 633 citations and The Cureus Journal of Medical Science with 448 citations highlights the academia interest in the use of AI, genAI and LLMs in healthcare
# Sources Articles
1 ARXIV 10,237
2 ADV NEUR IN 1,021
3 NATURE 980
4 J MED INTERNET RES 633
5 IEEE ACCESS 618
6 LECT NOTES COMPUT SC 464
7 COMPUT HUM BEHAV 454
8 CUREUS J MED SCIENCE 448
9 SCIENCE 448
10 INT J INFORM MANAGE 405
Table 2. Most Local Cited Sources
co-citation analysis, bibliographic coupling, keyword co-occurrence and other techniques
4.5. Most Relevant Countries
The Table 3 shows the country distribution of research articles based on corresponding authors, distinguishing between single-country publications (SCP) and multi-country publications (MCP). The USA leads with 610 articles (30.4%), of which 18% involve international collaboration (MCP). China (10.5%) follows, but with a higher proportion of MCPs (33.3%), indicating a stronger commitment to international research. The United Kingdom (7.4%) and Canada (2.9%) have some of the highest MCP percentages (45.9% and 41.4% respectively), reflecting their active participation in global collaborations. Other notable contributors include Germany (5.1%), Australia (4.2%) and India (3.2%), all of which also have a significant share of international partnerships, indicating the increasing globalisation of AI and LLM research. For better visualization, we also provide Figue 3.
This highlights the trend that the US and China are the leading forces driving innovation in AI and LLM research, reflecting their dominant roles in technological advancements and AI development, with technologies such as OpenAI’s ChatGPT in USA and DeepSeek in China. Meanwhile, the European Union (EU) tends to focus on regulation, as seen in policies such as the GDPR and the EU AI Act, emphasising ethical oversight and legal frameworks rather than direct competition in AI innovation.
# Country Articles Articles % SCP MCP MCP %
1 USA 610 30.4 500 110 18.0
2 China 210 10.5 140 70 33.3
3 United Kingdom 148 7.4 80 68 45.9
4 Germany 102 5.1 69 33 32.4
5 Australia 85 4.2 54 31 36.5
6 India 65 3.2 42 23 35.4
7 Canada 58 2.9 34 24 41.4
8 Spain 57 2.8 38 19 33.3
9 Italy 56 2.8 41 15 26.8
10 Netherlands 33 1.6 19 14 42.4
Table 3. Country Contribution Based on Corresponding Authors
Figure 3. Country Contribution Based on Corresponding Authors. Note: MCP represents multi-country publications, while SCP indicates the ratio of single-country publications.
4.6. Thematic Evolution
@@
The thematic evolution of AI ethics and trustworthiness in LLM concerns is illustrated in the Figure 4, which maps the shifts in research focus from 2019 to 2025. Initially, between 2019 and 2022, research focused predominantly on broad terms such as artificial intelligence, ethics, machine learning, language models and AI – reflecting general concerns about AI development and ethical challenges.

### v2

@@
Abstract
The rapid proliferation of Large Language Models (LLMs) has raised significant trustworthiness and ethical concerns. Despite the widespread adoption of LLMs across domains, there is still no clear consensus on how to define and operationalise trustworthiness. This study aims to bridge the gap between theoretical discussion and practical implementation by analysing research trends, definitions of trustworthiness, and practical techniques. We conducted a bibliometric mapping analysis of 2,006 publications from Web of Science (2019-2025) using the Bibliometrix, and manually reviewed 68 papers. We found a shift from traditional AI ethics discussion to LLM trustworthiness frameworks. We identified 18 different definitions of trust/trustworthiness, with transparency, explainability and reliability emerging as the most common dimensions. We identified 20 strategies to enhance LLM trustworthiness, with fine-tuning and retrieval-augmented generation (RAG) being the most prominent. Most of the strategies are developer-driven and applied during the post-training phase. Several authors propose fragmented terminologies rather than unified frameworks, leading to the risks of “ethics washing,” where ethical discourse is adopted without a genuine regulatory commitment. Our findings highlight: persistent gaps between theoretical taxonomies and practical implementation, the crucial role of the developer in operationalising trust, and call for standardised frameworks and stronger regulatory measures to enable trustworthy and ethical deployment of LLMs.
Index Terms:
@@ Abstract
I Introduction
@@
Artificial Intelligence (AI) has become a cornerstone of technological progress, yet remains at the centre of ethical debate. More than 200 ethical guidelines and principles have been proposed to guide its development and use, reflecting the growing recognition of its societal impact [1, 2]. These guidelines, proposed by public organisations, society, academia and private companies, do not provide objective guidance to developers, but rather abstract principles that are not practical for developers to operationalise, e.g, transparency, privacy, justice & fairness [3, 4, 5].
@@
As an evolution of AI system that employs advanced deep learning techniques, Large Language Models (LLMs) are trained on vast datasets and capable of generating novel text based on the data it was trained on [6]. Their rapid adoption has sparked both enthusiasm and concern, as their potential for negative outcomes – e.g., spreading misinformation, perpetuating biases, hallucination – has become evident [7, 8].
@@
The swift popularization of LLMs have encouraged researchers in understanding the specific ethical dimensions of this new generative AI (genAI), introducing them mainly as trustworthiness taxonomies, i.e., a rebranded set of similar AI ethical principles [6, 7, 8]. Trustworthiness in LLMs can be characterized in terms of reliability, safety, fairness, resistance to misuse, explainability & reasoning, social norm and robustness [7]. Thus, the focus of AI ethics discussions has moved to the trustworthiness of LLM systems. The current debate on LLM trustworthiness largely focus on providing definitions, benchmarks and evaluation methods [6, 7, 8].
@@
Analogously with AI ethics, the proliferation of trustworthiness taxonomies for LLM has not led to a clear consensus on definitions or a practical framework to guide the development of these systems [9]. Despite the growing body of literature on LLM trustworthiness, there is still no unified set of practical guidelines to help developers of LLM-based systems mitigate risks throughout the development and deployment process. This gap between theory and practice not only hinders the creation of more trustworthy LLM systems but also masks deeper structural problems that have persisted since the early days of AI ethics. While only a limited number of actors develop foundational LLMs, the widespread integration of these models into downstream systems highlights the urgent need for trustworthiness-enhancing practices that can be applied by practitioners leveraging on LLMs. This study aims to bridge the gap between theory and practice by investigating key research trends, definitions of LLM trustworthiness, and practical means of operationalising it found in the literature.
We conducted a bibliometric mapping of trust, trustworthiness, and ethics in LLMs using the Web of Science database. An initial set of 2,006 studies published between 2019 and 2025 was collected and analyzed to identify influential authors, key research trends, and emerging approaches. To deepen our analysis, we filtered a subset of 68 papers that explicitly mentioned “trust*” in their abstracts. These papers were manually reviewed to identify definitions and practical strategies that developers can use to enhance trustworthiness across the LLM development lifecycle. By bringing these perspectives together, we aim to provide both a theoretical and practical basis for promoting more ethical and trustworthy LLMs.
@@
  • AI ethics research focus has shifted from abstract guiding principles to abstract LLM trustworthiness taxonomies. Several authors propose fragmented terminologies rather than unified frameworks, leading to the risks of “ethics washing,” where ethical discourse is adopted without a genuine regulatory commitment.
  • We compiled 18 definitions of trust / trustworthiness in LLMs proposed by different authors. Transparency, explainability and reliability are the most frequent terms associated. Most of reviewed papers build on Mayer et al.’s [10] definition of organizational described as a three factor construct: ability, capacity to perform tasks effectively; benevolence, intent to act in users’ best interest; integrity, adherence to ethical stands.
  • We identified 20 strategies for enhancing trustworthiness in LLMs. Most strategies occur during the post-training phase and are developer-driven, with supervised fine-tuning, retrieval-augmented generation (RAG), and knowledge graphs being the most cited techniques. Developers are responsible for the majority of these strategies (16 out of 20), highlighting an imbalance in accountability and underscoring their central role in operationalising trustworthiness.
II Related Work
Several authors have analysed AI ethical guidelines, proposing novel sets of principles, such as [11, 2, 1, 12]. Correa et al. [2] performed a meta-analysis of 200 AI ethics documents, summarising them in 17 principles, where 5 principles emerge as more recurrent: 1) Transparency / Explainability / Auditability, 2) Reliability / Safety / Security / Trustworthiness, 3) Justice / Equity / Fairness/ Non-discrimination, 4) Privacy and 5) Accountability / Liability. Morley et al. [13] provide an overview of 106 publicly available tools designed to help operationalize ethical principles in AI. Based on five principles – beneficence, non-maleficence, autonomy, justice, and explicability – the authors map practical techniques to the machine learning lifecycle phase , illustrating how these tools can support ethical AI development in practice. Studies on AI ethics that emerged before the rise of LLMs and generative AI offer a useful starting point for exploring trustworthiness in these technologies. They provide the background and foundational frameworks on which future research into LLM trustworthiness can build, since LLMs are ultimately part of the broader field of AI.
@@ II Related Work
Regarding LLM and genAI trust/trustworthiness, recent studies are proposing similar set of theoretical taxonomies along with evaluation methods against such taxonomies. Liu et al. [7] propose a taxonomy for evaluating trustworthiness in LLM in seven major categeories: reliability, safety, fairness, resistance to misuse, explainability and reasoning, adherence to social norms, and robustness. Although the paper is not a literature review, it does describe each of the major categories with examples. However, the authors fail to provide a clear definition of what is trustworthy and do not provide practical guidance for developers on how to operationalise their taxonomy. Sun et al. [6] and Wang et al. [8] propose different sets of principles for trustworthiness in LLM, such as truthfulness, safety, fairness, robustness, privacy, machine ethics, transparency, accountability. Nonetheless, they do not provide developers with a practical guidance on how to operationalise these principles.
@@
Dhoska and Bebi [14] performed a bibliometric analysis of trustworthy in AI in the Scopus database, showing e.g., relevant authors, affiliations, countries and most cited papers. Nevertheless, their study focused on AI rather than LLM specifically, and they did not present the importance of trustworthiness in AI in their discussion, focusing only on the results of the bibliometric metrics and not on the content of the papers analysed. Albahri et al. [15] conducted a systematic science mapping analysis in the context of trustworthy and explainable AI (XAI) in healthcare systems, and found various XAI techniques for developers to operationalise. Unlike their approach, we aim to provide means to operationalise more of the trustworthiness categories.
Despite the aforementioned efforts in this area, our work differs by exploring not only a bibliometric analysis and definitions of trustworthiness in LLM, but also practical means for developers to improve LLM trustworthiness throughout the LLM lifecycle.
III Methodology
@@
Bibliometric analysis is a quantitative research method used to identify trends, influential works, and key research themes in a given domain [16]. To achieve this, we utilize Bibliometrix, an open-source R package specifically designed for bibliometric and scientometric studies [17]. We extracted data from Web of Science (WoS) Core Collection during 21st of February 2025, as this database contain high-quality, peer-reviewed publications in AI and Software Engineering, and provides high quality of publications and reliability in its indexing of high-ranking journals [18]. The search query was formulated using a combination of relevant broad keywords seeking for a comprehensive dataset:
TS=((“Large Language Model*” OR LLM OR “Generative AI” OR “genAI”) AND (Trust* OR Ethic*))
@@
Bibliometric analysis: After filtering out duplicates, irrelevant papers and LLM as an author abbreviation, we obtained a final dataset of 2,006 publications. This number of studies discussing LLMs is remarkable as we reduced the year range of our search to only 6 years due to the fact that LLM is a new technology. A BibTex file was extracted from WoS and imported into Bibliometrix, where we carried out the bibliometric analysis to answer RQ 1.
@@
Further manual reading: After exploring the broader set with Bibliometrix, we conducted a manual reading of a reduced dataset. In order to obtain more insightful information related to our research focus, we filtered papers that had “trust*” in the abstract. Thus, a subset of 68 papers was obtained, and we used this subset to answer RQs 2 and 3.
IV Results
@@
This study examines the landscape of trust, trustworthiness and ethics research in LLM and genAI from 2019 to 2025. A total of 2,006 relevant studies published in 1,160 different venues over the past six years were analysed. These publications were authored by 7,901 researchers, with an average citation count of 8.901 per document. Notably, collaborative research dominates the field, with 96.6% of the studies involving multiple authors (7,632 in total), while only 3.4% were produced by single authors (269 individuals).
@@
The study shows a sharp rise in publications on LLM trustworthiness and ethics from 2019 to early 2025, growing from 3 articles in 2019 to 1,434 in 2024, as shown in Figure 1. By February 2025, 213 articles had already been published, indicating continued momentum ( 1,278 projected by year-end). Collaborative research dominated, with 96.6% of studies involving multiple authors.
Figure 1: Evolution of Studies Per Year - 2019 - 2025.
Figure 2 shows the top articles ranked by total number of citations, without any filtering. The two most globally cited articles are “ChatGPT for good? On opportunities and challenges of large language models for education” by Kasneci et al. [19] and “Opinion Paper: “So what if ChatGPT wrote it?” Multidisciplinary perspectives on opportunities, challenges and implications of generative conversational AI for research, practice and policy” by Dwivedi et al. [20], both with 1,135 global citations. The third most cited article “ChatGPT Utility in Healthcare Education, Research, and Practice: Systematic Review on the Promising Perspectives and Valid Concerns” by Sallam Malik [21], with 838 global citations.
Figure 2: Most Globally Cited Articles (2022–2024).
Figure 3: Thematic Evolution: 2019-2025.
@@
Thematic evolution of AI ethics, trustworthiness, and LLMs is shown in Figure 3, which maps the shifts in research focus from 2019 to 2025. From 2019 to 2022, research primarily addressed broad topics such as artificial intelligence, ethics, machine learning, and language models, reflecting general concerns over AI development and ethical challenges.